OpenStudy (anonymous):

two kinds of nuts are mixed to sell at 70 cents per bag. one kind sells at 50 cents per bag; the other at 80 cents per bag. How much of each kind should be used in 300 bags of the mixture? (use 2 variables to solve)

OpenStudy (anonymous):

@amistre64 heree's what i dod. x= amt of bagsfor 1st kind y= amt of bags for 2nd kind 50x+20y= 70(x+y) x+y=300

OpenStudy (anonymous):

do i use number of nuts or bags? o.o

OpenStudy (anonymous):

i think i got first equation right, but i don't wknow what to set it equal to

OpenStudy (anonymous):

oh wait, i need a variable for total amt.

OpenStudy (amistre64):

two kinds of nuts are mixed to sell at 70 cents per bag. one kind sells at 50 cents per bag; 50x the other at 80 cents per bag. 80y How much of each kind should be used in 300 bags of the mixture? x+y = 300 so: 50x + 80y = 70(300) does this seem right?

OpenStudy (anonymous):

ahh it makes sense now with the 70 cents per bag times 300. ty once again!

OpenStudy (amistre64):

youre welcome